swarnbhashm is a cosmetic medicene in which found 400ppm colydee gold.what is the mass% of gold?
1
Expert's answer
2017-06-30T04:37:06-0400
Parts per million is the mass ratio between the pollutant component and the solution - and ppm is defined as
ppm = 1000000 m_c/m_s
where
mc = mass of component (kg)
ms = mass of solution (kg)

Percent concentration by mass is defined as the mass of solute divided by the total mass of the solution and multiplied by 100%. So,
c %=m_solute/m_solution ∙100%
Thus,
c %=ppm/1000000∙100%=400/1000000∙100%=400/10000=0.04%

Answer: 0.04%